I forgot how to complete this equation;
Equation
A.)2(3)2-3(5)
B.)3(3)2-2(5)
The answer I have for A. is 31, for B. the answer I have is 71. I subtracted them and got 40. But 40 is not an answer choose????

Answer by JBarnum(2146)

theres no equation present...for there to be an equation there needs to be an "Equal" sign
why would u subtract one from the other?
Where is the rest of the problem?
all i see is A=12-15=-3 and B=18-10=28
2*3*2-3*5=-3
3*3*2-2*5=28
So theres obviously something missing with this problem.
Please write clearly, explain what the equation is suppose to looklike if you cant duplicate the symbols. put the entire question in a post then put what you have tried to do.
We arent Mind Readers.